# Stable Offline Value Function Learning with Bisimulation-based Representations

### v2

@@ 1 Introduction
In seeking such representations, we turn to $\pi$-bisimulation algorithms. These algorithms define a metric to capture behavioral similarity between state-action pairs such that similarity is based on immediate rewards received and the similarity of next state-action pairs visited by $\pi$ (Castro, 2019). The algorithms then use this metric to learn representations such that state-action pairs that are similar under this metric have similar representations. Ultimately, the goal of $\pi$-bisimulation methods is to learn representations such that state-actions pairs with similar values under $\pi$ also have similar representations (see Figure 1). While these algorithms have shown promise in improving the expected return of rl algorithms, it remains unclear whether they contribute to stability (Castro et al., 2023; Zhang et al., 2021; Castro et al., 2022). In this paper, we aim to understand whether $\pi$-bisimulation-based representations stabilize value function learning.
Figure 1: The figure illustrates the native state-action representations $\mathcal{X}$ and $\pi$-bisimulation representations $\Phi_{\pi-\textsc{bisim}}$. $\pi$-bisimulation algorithms use a similarity function $k$ that captures similarity between state-action pairs based on immediate rewards and similarity of next state-action pairs under $\pi$ to shape their representations. Ultimately, the goal of $\pi$-bisimulation methods is to learn representations such that state-actions pairs with similar values under $\pi$ also have similar representations. The function $k$ outputs high values within the blue (and orange) state-actions but low values between blue and orange state-actions. Therefore, the blue (and orange) state-actions have similar representations, but different representations between the distinct colors.
In this work, we focus on offline value function learning. Given a fixed, offline dataset generated by unknown and possibly multiple behavior policies, the goal is to estimate the value function of a fixed, target policy. We introduce kernel representations for offline policy evaluation (krope), a bisimulation-based representation learning algorithm. krope defines a kernel that captures similarity between state-action pairs based on immediate rewards received and similarity of next state-action pairs under the target policy. It then shapes the state-action representations such that state-action pairs that are similar according to this kernel have similar representations. We use krope as the representative algorithm for the class of bisimulation-based representation learning algorithms to investigate the following question:
@@
    We empirically validate that krope representations lead to more stable and accurate offline value function learning compared to non-bisimulation baselines (Section 4).
@@ 2.2 Offline Policy Evaluation and Value Function Learning
where the expectations are taken with respect to the randomness of the dataset $\mathcal{D}$ and ${\pi_{e}}$. Note that $\mathbb{E}[\Phi^{\top}\Phi]$ is the feature covariance matrix. Assuming lspe converges, it will converge to the same fixed-point as td(0) (Szepesvari, 2010), which we denote as $\theta_{\text{{lspe}}}$. In this work, we follow a two-stage approach to applying lspe: we first obtain the encoder $\phi$ either through representation learning or using the native features of the mdp, and then feed the obtained $\phi$ along with $\mathcal{D}$ and ${\pi_{e}}$ as input to lspe, which outputs $\hat{q}_{\theta}^{{\pi_{e}}}$ (Nedic & Bertsekas, 2003; Chang et al., 2022). This two-stage approach of learning a linear function on top of fixed representations is called the linear evaluation protocol (Chang et al., 2022; Farebrother et al., 2023; 2024; Grill et al., 2020; He et al., 2020). This protocol enables us to cleanly analyze the nature of the learned representations within the context of well-understood value function learning algorithms such as lspe. In Appendix A, we include the pseudo-code for lspe.
2.3 Stable, Realizable, and Generalizable Representations
We define stability of lspe and related td-methods following Ghosh & Bellemare (2020):
@@ 2.3 Stable, Realizable, and Generalizable Representations
When determining the stability of lspe, we have following proposition from prior work:
Proposition 1 (Asadi et al. (2024); Wang et al. (2021a)).
@@
In addition to stability, we also care about the realizability and generalizability of $\Phi$. We say $\Phi$ is a realizable representation if $q^{{\pi_{e}}}\in\text{Span}(\Phi),$ where $\text{Span}(\Phi)$ is the subspace of all expressible action-value functions with $\Phi$. Note that even if $\Phi$ is a realizable and stable representation, lspe may not recover the $q^{{\pi_{e}}}$ solution (Sutton & Barto, 2018). While generalization can have multiple interpretations, we say $\Phi$ generalizes well if the state-action features that are close in the representation space also have similar $q^{\pi_{e}}$ values (Lyle et al., 2022; Lan et al., 2021).
@@ 2.4 Related Works
Recently, there has been lot of interest in $\pi$-bisimulation algorithms for better generalization (Ferns et al., 2004; 2011; Ferns & Precup, 2014; Castro, 2019; Zang et al., 2023b). These algorithms measure similarity between two state-action pairs based on immediate rewards received and the similarity of next state-action pairs visited by $\pi$. These algorithms first define a distance metric that captures this $\pi$-bisimilarity, and then use this metric to learn representations such that $\pi$-bisimular states have similar representations (Castro et al., 2022; Castro, 2019; Zhang et al., 2021; Castro et al., 2023; Chen & Pan, 2022; Kemertas & Jepson, 2022). Zhang et al. (2021); Castro (2019) introduced a $\pi$-bisimulation learning algorithm but assume that the transition dynamics are either deterministic or Gaussian. Gelada et al. (2019) introduced a method closely related to bisimulation methods but required a reconstruction loss to work in practice. Castro et al. (2022) introduced mico which allows for stochastic transition dynamics and no reconstruction loss, but was difficult to theoretically analyze. To overcome this difficulty, Castro et al. (2023) took a kernel perspective of $\pi$-bisimulation methods, which made their algorithm amenable to theoretical analysis. To the best of our knowledge, no works have studied the stability properties of $\pi$-bisimulation algorithms. In our work, we address this gap in the literature. We first extend Castro et al. (2023)’s kernel-based formulation from states to state-actions, and then show that this formulation stabilizes offline value function learning. The proofs for krope’s basic theoretical properties (Section 3.1) follow those by Castro et al. (2023). Our stability-related theoretical results (Sections 3.2 and 3.3) and empirical analysis (Section 4) are novel to this work.
@@ 3 Kernel Representations for Offline Policy Evaluation
We now present our bisimulation-based representation learning algorithm, kernel representations for ope (krope). We present the desired krope kernel, define the krope operator, present its theoretical properties, prove stability properties of krope representations, and present a practical learning algorithm to learn them. We defer the proofs to Appendix B.
3.1 krope Kernel and Operator
@@
Prior $\pi$-bisimulation works define similarity between states in terms of the immediate rewards received and similarity of next states under $\pi$ (see Figure 1) (Castro, 2019). In this work, we follow Castro et al. (2023) and define a kernel $k^{{\pi_{e}}}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ that captures this notion of similarity under ${\pi_{e}}$, but for pairs of state-actions. We refer to $k^{\pi_{e}}$ as the krope kernel.
$k^{{\pi_{e}}}(s_{1},a_{1};s_{2},a_{2})=k_{1}(s_{1},a_{1};s_{2},a_{2})+\gamma k_{2}(k^{{\pi_{e}}})(P^{\pi_{e}}(\cdot|s_{1},a_{1}),P^{\pi_{e}}(\cdot|s_{2},a_{2})].$ (2)
@@
where $k_{1}(s_{1},a_{1};s_{2},a_{2}):=1-\frac{|r(s_{1},a_{1})-r(s_{2},a_{2})|}{\left|r_{\text{max}}-r_{\text{min}}\right|}$ and $k_{2}(k^{\pi_{e}})(P^{\pi_{e}}(\cdot|s_{1},a_{1}),P^{\pi_{e}}(\cdot|s_{2},a_{2})):=\mathbb{E}_{s_{1}^{\prime},a_{1}^{\prime}\sim P^{\pi_{e}}(\cdot|s_{1},a_{1}),s_{2}^{\prime},a_{2}^{\prime}\sim P^{\pi_{e}}(\cdot|s_{2},a_{2})}[k^{\pi_{e}}(s_{1}^{\prime},a_{1}^{\prime};s_{2}^{\prime},a_{2}^{\prime})]$. Here, $k_{1}$ measures short-term similarity based on rewards received and $k_{2}$ measures long-term similarity between probability distributions by measuring similarity between samples of the distributions according to $k^{\pi_{e}}$ (Castro et al., 2023).
Given this definition of the krope kernel, we now present an operator that converges to $k^{\pi_{e}}$:
Definition 2 (krope operator).
[…]
Proposition 2, proved in Appendix B.1, tells us that for some initial kernel $k$, repeatedly applying $\mathcal{F}^{\pi_{e}}$ to it will result in convergence $k^{{\pi_{e}}}$. Ultimately, $k^{\pi_{e}}$ outputs a high (or low) similarity measure for two state-action pairs if their action-values under ${\pi_{e}}$ are similar (or dissimilar). This intuition is formalized in Lemma 3, which states that the absolute action-value difference between any two state-action pairs under ${\pi_{e}}$ is upper-bounded by the distance function induced by $k^{\pi_{e}}$ plus an additive constant. Since krope’s contraction, metric space completeness, and fixed-point uniqueness properties are similar to Castro et al. (2023)’s kernel, we defer the details to Appendix B.1.
@@ 3.2 Stability of krope Representations
Theorem 1, proved in Appendix B.2, tells us that krope representations stabilize ope with lspe. Intuitively, they are stable since when $(\mathbb{E}[\Phi^{\top}\Phi]))^{-1}\mathbb{E}[\gamma\Phi^{\top}P^{{\pi_{e}}}\Phi]$’s spectral radius is less than $1$, each update to $\theta_{t}$ in Equation (1) is non-expansive. That is, each update brings $\theta_{t}$ closer to $\theta_{\textsc{lspe}}$.
@@ 3.3 Connection to Bellman Completeness
Let the induced distance function due to $k^{\pi_{e}}$ be
@@
$\mathcal{L}_{\text{{krope}}}(\omega):=\mathbb{E}_{\mathcal{D}}\Biggl{[}\biggl{(}\underbrace{1-\frac{\left|r(s_{1},a_{1})-r(s_{2},a_{2})\right|}{\left|r_{\text{max}}-r_{\text{min}}\right|}+\gamma\mathbb{E}_{{\pi_{e}}}[\tilde{k}_{\bar{\omega}}(s_{1}^{\prime},a_{1}^{\prime};s_{2}^{\prime},a_{2}^{\prime})]}_{\text{target estimate}}-\underbrace{\vphantom{\bigg{|}}\tilde{k}_{\omega}(s_{1},a_{1};s_{2},a_{2})}_{\text{current estimate}}\biggr{)}^{2}\Biggr{]},$ (5)
where the state-action pairs are sampled from $\mathcal{D}$, and $\bar{\omega}$ are weights of the target network that are periodically copied from $\omega$ (Mnih et al., 2015). In this work, we use krope as an auxiliary task, which introduces only a learning rate as an additional hyperparameter. We note that this fixed-point optimization procedure is similar to how the action-value function is learned in other rl fixed-point algorithms such as fitted q-evaluation (fqe) (Le et al., 2019).
@@ 4 Empirical Results
In this section, we present our empirical study designed to answer the following questions.
  1. 1.
    Does krope lead to stable representations with good realization and generalization?
  2. 2.
    Do krope representations lead to stable msve and low msve?
@@ 4.1 Empirical Setup
We consider the following four non-bisimulation auxiliary representation learning algorithms that are typically paired with fqe for stability: 1) krope (ours), 2) bcrl-exp-na, which simultaneously optimizes three objectives: immediate reward prediction, next-state self-prediction loss, and maximization of the exploratory nature of $\mathbb{E}[\Phi^{\top}\Phi]$ by improving its condition number (Chang et al., 2022), 3) Absolute dr3 regularizer (Kumar et al., 2021; Ma et al., 2024), which promotes stability by minimizing the feature co-adaptation between successive features, i.e., $\phi(s,a)^{\top}\phi(s^{\prime},a^{\prime})$, and 4) beer regularizer (He et al., 2024), which is similar to dr3 but lower bounds $\phi(s,a)^{\top}\phi(s^{\prime},a^{\prime})$. In all cases, the penultimate layer features of fqe’s action-value encoder $\phi_{\xi^{\prime}}$ are fed into lspe for ope. Our experiments focus on analyzing the properties of $\phi_{\xi^{\prime}}$. Note that since bcrl was not designed as an auxiliary task (Chang et al., 2022), we evaluate it as a non-auxiliary (na) task algorithm. We provide additional details on the baselines in Appendix C.
@@
We conduct our evaluation on a variety of domains: 1) Garnet mdps, which are a class of tabular stochastic mdps that are randomly generated given a fixed number of states and actions (Archibald et al., 1995); 2) $4$ dm Control environments: CartPoleSwingUp, CheetahRun, FingerEasy, WalkerStand (Tassa et al., 2018); and 3) $9$ d4rl datasets (Fu et al., 2020; 2021). The first domain enables us to analyze the algorithms’ performance across a wide range of stochastic tabular mdps. The second and third set of domains test the algorithms in continuous higher-dimensional state-action environments. Due to space constraints, we defer the d4rl results to Appendix C.3.2.
4.2 Analyzing Fundamental Properties of the Learned Representations
In this set of experiments on the Garnet mdps domain, we answer our first question of whether krope representations lead to stable representations with good realization and generalization properties. We present the results in Figure 2. Our Garnet mdps were generated with $8$ states and $5$ actions, with a total of $|\mathcal{X}|=40$ state-actions, and each native $(s,a)$ representation is a $1$-hot vector. In these experiments, the native representation is fed into a linear encoder with a bias component and no activation function. All algorithms are trained for $500$ epochs and we report the results by evaluating the final learned representations for different latent dimensions $d$.

### v3

@@ 1 Introduction
Figure 1: The figure illustrates the native state-action representations $\mathcal{X}$ and $\pi$-bisimulation representations $\Phi_{\pi-\textsc{bisim}}$. $\pi$-bisimulation algorithms use a distance function $d$ that captures differences between state-action pairs based on immediate rewards and differences of next state-action pairs under $\pi$ to shape their representations. Ultimately, the goal of $\pi$-bisimulation methods is to learn representations such that state-actions pairs with similar action-values under $\pi$ also have similar representations. The function $d$ outputs low values within the blue (and orange) state-actions but high values between blue and orange state-actions. Therefore, the blue (and orange) state-actions have similar representations, but different representations between the distinct colors.
In seeking such representations, we turn to $\pi$-bisimulation algorithms. These algorithms define a metric to capture behavioral similarity between state-action pairs such that similarity is based on immediate rewards received and the similarity of next state-action pairs visited by $\pi$ (Castro, 2019). The algorithms then use this metric to learn representations such that state-action pairs that are similar under this metric have similar representations. Ultimately, the goal of $\pi$-bisimulation methods is to learn representations such that state-actions pairs with similar values under $\pi$ also have similar representations (see Figure 1). While these algorithms have shown promise in improving the expected return of rl algorithms, it remains unclear whether they contribute to stability (Castro et al., 2023; Zhang et al., 2021; Castro et al., 2022). In this paper, we aim to understand whether the $\pi$-bisimulation-based representations stabilize value function learning.
In this work, we focus on offline value function learning. Given a fixed, offline dataset generated by unknown and possibly multiple behavior policies, the goal is to estimate the value function of a fixed, target policy. Towards establishing the stability properties of $\pi$-bisimulation representations, we introduce kernel representations for offline policy evaluation (krope). krope defines a kernel that captures similarity between state-action pairs based on immediate rewards received and similarity of next state-action pairs under the target policy. It then shapes the state-action representations such that state-action pairs that are similar according to this kernel have similar representations. We use krope as the representative algorithm for the class of bisimulation-based representation learning algorithms to investigate the following question:
@@
    We empirically validate that krope representations lead to more stable and accurate offline value function learning compared to baselines (Section 4).
@@ 2.2 Offline Policy Evaluation and Value Function Learning
where the expectations are taken with respect to the randomness of the dataset $\mathcal{D}$ and ${\pi_{e}}$. Note that $\mathbb{E}[\Phi^{\top}\Phi]$ is the feature covariance matrix. Assuming lspe converges, it will converge to the same fixed-point as td(0) (Szepesvari, 2010), which we denote as $\theta_{\text{{lspe}}}$. In this work, we follow a two-stage approach to applying lspe: we first obtain the encoder $\phi$ either through representation learning or using the native features of the mdp, and then feed the obtained $\phi$ along with $\mathcal{D}$ and ${\pi_{e}}$ as input to lspe, which outputs $\hat{q}_{\theta}^{{\pi_{e}}}$ (Nedic & Bertsekas, 2003; Chang et al., 2022). This two-stage approach of learning a linear function on top of fixed representations is called the linear evaluation protocol (Chang et al., 2022; Farebrother et al., 2023, 2024; Grill et al., 2020; He et al., 2020). This protocol enables us to cleanly analyze the nature of the learned representations within the context of well-understood value function learning algorithms such as lspe. In Appendix A, we include the pseudo-code for lspe.
2.3 Stable Representations and $q^{\pi_{e}}$-Consistency
We define stability of lspe and related td-methods following (Ghosh & Bellemare, 2020):
@@ 2.3 Stable Representations and $q^{\pi_{e}}$-Consistency
When determining the stability of lspe, we have the following proposition from prior work:
Proposition 1 (Asadi et al. 2024; Wang et al. 2021a).
@@
In addition to stability, we also want the state-action features to be such that state-action features that are close in the representation space also have similar $q^{\pi_{e}}$ values (Lyle et al., 2022; Lan et al., 2021). In this work, we call this property $q^{\pi_{e}}$-consistency since the learned representations are consistent with the $q^{\pi_{e}}$ values.
@@ 2.4 Related Works
Recently, there has been lot of interest in $\pi$-bisimulation algorithms for better generalization (Ferns et al., 2004, 2011; Ferns & Precup, 2014; Castro, 2019; Zang et al., 2023b). These algorithms measure similarity between two state-action pairs based on immediate rewards received and the similarity of next state-action pairs visited by $\pi$. These algorithms first define a distance metric that captures this $\pi$-bisimilarity, and then use this metric to learn representations such that $\pi$-bisimular states have similar representations (Castro et al., 2022; Castro, 2019; Zhang et al., 2021; Castro et al., 2023; Chen & Pan, 2022; Kemertas & Jepson, 2022). Castro (2019) introduced a $\pi$-bisimulation learning algorithm but assumed that the transition dynamics are deterministic. Zhang et al. (2021) introduced a bisimulation algorithm that allowed for Gaussian dynamics and demonstrated its effectiveness in discarding distracting features for control. Castro et al. (2022, 2023) introduced bisimulation-based algorithms for control that allow for stochastic transition dynamics. To the best of our knowledge, there is still a gap in the literature as to whether the general class of these learned $\pi$-bisimulation-based representations stabilize offline value function learning. To address this gap, we introduce krope, an algorithm that leverages Castro et al. (2023)’s kernel perspective on similarity metrics. We first explicitly define a kernel that captures the relationship between state-action features in latent space in terms of the $\pi$-bisimulation relation. This definition allows us to immediately establish stability since we can now analyze the spectral properties of offline value function learning algorithms. The proofs for krope’s basic theoretical properties (Section 3.1) follow those by Castro et al. (2023). Our stability-related theoretical results (Sections 3.2 and 3.3) and empirical analysis (Section 4) are novel to this work.
@@ 3 Kernel Representations for Offline Policy Evaluation
We now present our bisimulation-based representation learning algorithm, kernel representations for ope (krope). We present the desired krope kernel, prove stability properties of krope representations, and present a practical learning algorithm to learn them. We defer the proofs to Appendix B.
3.1 krope Kernel
@@
Prior works typically define a distance metric to capture the notion of $\pi$-bisimilarity between two states: the distance between states is in terms of the immediate rewards received and differences of next states under $\pi$ (see Figure 1) (Castro, 2019). In this work, we follow Castro et al. (2023) and take a perspective of kernels instead of distances. We first define a kernel that captures $\pi$-bisimilarity $k^{{\pi_{e}}}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, but for pairs of state-actions and under ${\pi_{e}}$:
$k^{{\pi_{e}}}(s_{1},a_{1};s_{2},a_{2})=k_{1}(s_{1},a_{1};s_{2},a_{2})\\ +\gamma k_{2}(k^{{\pi_{e}}})(P^{\pi_{e}}(\cdot|s_{1},a_{1}),P^{\pi_{e}}(\cdot|s_{2},a_{2})].$ (2)
@@
where $k_{1}(s_{1},a_{1};s_{2},a_{2}):=1-\frac{|r(s_{1},a_{1})-r(s_{2},a_{2})|}{\left|r_{\text{max}}-r_{\text{min}}\right|}$ and $k_{2}(k^{\pi_{e}})(P^{\pi_{e}}(\cdot|s_{1},a_{1}),P^{\pi_{e}}(\cdot|s_{2},a_{2})):=\mathbb{E}_{s_{1}^{\prime},a_{1}^{\prime}\sim P^{\pi_{e}}(\cdot|s_{1},a_{1}),s_{2}^{\prime},a_{2}^{\prime}\sim P^{\pi_{e}}(\cdot|s_{2},a_{2})}[k^{\pi_{e}}(s_{1}^{\prime},a_{1}^{\prime};s_{2}^{\prime},a_{2}^{\prime})]$. Here, $k_{1}$ measures short-term similarity based on rewards received and $k_{2}$ measures long-term similarity between probability distributions by measuring similarity between samples of the distributions according to $k^{\pi_{e}}$ (Castro et al., 2023). We refer to $k^{\pi_{e}}$ as the krope kernel.
Remarks on $k^{\pi_{e}}$.
We now discuss the trade-offs of $k^{\pi_{e}}$. While $k^{\pi_{e}}$ is different from prior work, it benefits and suffers from the same advantages and disadvantages discussed in Castro et al. (2023).
By definition $k_{2}$ is a function of independently coupling $(s_{1}^{\prime},a_{1}^{\prime})$ and $(s_{2}^{\prime},a_{2}^{\prime})$ (Castro et al., 2022, 2023). The advantage of this independence is that it enables a practical algorithm and is flexible as it make no assumptions on the nature of the transition dynamics of $P^{\pi_{e}}$. While independent coupling does raise a complication with stochastic dynamics and/or stochastic policies since self-similarity may be lower than similarity between different state-actions, i.e., $k^{\pi_{e}}(s_{1},a_{1};s_{1},a_{1})<k^{\pi_{e}}(s_{1},a_{1};s_{2},a_{2})$ for some $(s_{1},a_{1}),(s_{2},a_{2})\in\mathcal{X}$, $k^{\pi_{e}}$ still induces a valid $\pi$-bisimulation-based metric. To see its validity, consider the induced distance function by $k^{\pi_{e}}$, i.e., $\forall x,y\in\mathcal{X},d_{\text{{krope}}}(x,y):=k^{{\pi_{e}}}(x,x)+k^{{\pi_{e}}}(y,y)-2k^{{\pi_{e}}}(x,y)$. The metric $d_{\text{{krope}}}$ satisfies continuity (see Lemma 3 in Appendix B.1), which is an important property for metrics to satisfy (Lan et al., 2021). Lemma 3 states that the absolute action-value difference between any two state-action pairs under ${\pi_{e}}$ is upper-bounded by $d_{\text{{krope}}}$ plus an additive constant, where the additive constant arises due to the independent coupling. Thus far, prior work has removed the additive constant by assuming deterministic or Gaussian dynamics (Castro, 2019; Zhang et al., 2021). However, these assumptions are practically more restrictive. While one might expect the existence of the additive constant to hurt performance, prior work and our work (see Section 4) show that independent coupling actually yields strong performance (Pavse & Hanna, 2023a; Castro et al., 2023). Furthermore, in Section 3.2 we show that this property does not affect the stability properties of the learned representations. An interesting future direction will be to develop practically feasible approaches to eliminate the additive constant.
Since Lemma 3 and the associated contraction, metric space completeness, and fixed-point uniqueness properties are similar to Castro et al. (2023)’s kernel, we defer these results to Appendix B.1.
@@ 3.2 Stability of krope Representations
By adopting the kernel perspective, Theorem 1, proved in Appendix B.2, tells us that $\pi$-bisimulation-based krope representations stabilize ope with lspe. Intuitively, they are stable since when $(\mathbb{E}[\Phi^{\top}\Phi]))^{-1}\mathbb{E}[\gamma\Phi^{\top}P^{{\pi_{e}}}\Phi]$’s spectral radius is less than $1$, each update to $\theta_{t}$ in Equation (1) is non-expansive. That is, each update brings $\theta_{t}$ closer to $\theta_{\textsc{lspe}}$.
@@ 3.3 Connection to Bellman Completeness
Recall, that the induced distance function due to $k^{\pi_{e}}$ is
@@
$\mathcal{L}_{\text{{krope}}}(\omega)\\ :=\mathbb{E}_{\mathcal{D}}\Biggl{[}\biggl{(}\mathcal{K}_{\bar{\omega}}(s_{1},a_{1};s_{2};a_{2})-\vphantom{\bigg{|}}\tilde{k}_{\omega}(s_{1},a_{1};s_{2},a_{2})\biggr{)}^{2}\Biggr{]},$ (4)
where $\mathcal{K}_{\bar{\omega}}(s_{1},a_{1};s_{2};a_{2})=1-\frac{\left|r(s_{1},a_{1})-r(s_{2},a_{2})\right|}{\left|r_{\text{max}}-r_{\text{min}}\right|}+\gamma\mathbb{E}_{{\pi_{e}}}[\tilde{k}_{\bar{\omega}}(s_{1}^{\prime},a_{1}^{\prime};s_{2}^{\prime},a_{2}^{\prime})]$, the state-action pairs are sampled from $\mathcal{D}$, and $\bar{\omega}$ are weights of the target network that are periodically copied from $\omega$ (Mnih et al., 2015). In this work, we use krope as an auxiliary task, which introduces only an auxiliary task weight as the additional hyperparameter.
@@ 4 Empirical Results
In this section, we present our empirical study designed to answer the following question: do krope representations lead to stable msve and low msve?
@@ 4.1 Empirical Setup
We consider the following three classes (and total $6$) auxiliary representation learning algorithms that are typically paired with fqe for stability: I) bisimulation-metric based, II) model-based, and III) co-adaptation penalty-based. In class I, we consider 1) krope (ours), 2) deep bisimulation for control (dbc) (Zhang et al., 2021), 3) representations for ope (rope) (Pavse & Hanna, 2023a; Castro et al., 2022); in class II, we consider bellman complete representation learning (bcrl-exp-na) (Chang et al., 2022); and in class III, we consider 1) absolute dr3 regularizer (Kumar et al., 2021; Ma et al., 2024) and 2) beer regularizer (He et al., 2024). In all cases, the penultimate layer features of fqe’s action-value encoder $\phi_{\xi^{\prime}}$ are fed into lspe for ope. Note that since bcrl was not designed as an auxiliary task (Chang et al., 2022), we evaluate it as a non-auxiliary (na) task algorithm. We provide additional details on the baselines in Appendix C.
@@
We conduct our evaluation on a variety of domains: 1) Garnet mdps, which are a class of tabular stochastic mdps that are randomly generated given a fixed number of states and actions (Archibald et al., 1995); 2) $4$ dm Control environments: CartPoleSwingUp, CheetahRun, FingerEasy, WalkerStand (Tassa et al., 2018); and 3) $9$ d4rl datasets (Fu et al., 2020, 2021). The first domain enables us to analyze the algorithms’ performance across a wide range of stochastic tabular mdps. The second and third set of domains test the algorithms in continuous higher-dimensional state-action environments.
4.2 Analyzing Stability and $q^{\pi_{e}}$-Consistency Properties
In this set of experiments on the Garnet mdps domain, we analyze the stability and $q^{\pi_{e}}$ consistency properties of the learned representations. We present the results in Figure 2. Our Garnet mdps were generated with $8$ states and $5$ actions, with a total of $|\mathcal{X}|=40$ state-actions, and each native $(s,a)$ representation is a $1$-hot vector. In these experiments, the native representation is fed into a linear encoder with a bias component and no activation function. All algorithms are trained for $500$ epochs and we report the results by evaluating the final learned representations for different latent dimensions $d$.